# A practical identifiability criterion leveraging weak-form parameter estimation

### v2

@@ 3.1 Weak-form parameter estimation for systems with unobserved variables
In general, we convert between an equation in the form $\dot{U}=\Theta(U,W)$, with observed states, $y(t)=\Omega(U,W)$, to an input-output equation of the form $0=F(y)$. This process involves, first, ranking the model variables and their derivatives with the observed state variable at the lowest rank. Next, differential polynomials generated from the model equations are reduced to their characteristic set according to the new ranking using a differential elimination algorithm. After the characteristic set is obtained, the lowest ranking differential polynomial(s) will be the input-output equation(s).
@@
If the right-hand side of the input-output equation can be written in the form
$F(y)=\sum_{i=0}^{n}\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}H_{i}(y),$ (1)
@@
then it is clear that by convolving the input-output equation with a test function $\phi$ and applying integration by parts that we can generate a weak-form input-output equation of the form
$0=\sum_{i=0}^{n}\int_{0}^{T}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi(t))H_{i}(y)\textsf{d}t.$ (2)
However, not all nonlinear systems can be written in the form of equation (1). In the case that the right-hand side of the input-ouput equation takes the form,
$F(y)=\sum_{i=0}^{n}K_{i}(y)\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}H_{i}(y),$ (3)
@@
either substitutions of variables, estimation of higher order derivatives from data, or both will be required to compute terms that are not integrable by parts.
@@
Recently, our group has developed the Weak form Estimation of Nonlinear Dynamics (WENDy) method for parameter estimation [26]. Briefly, in the WENDy method, a system is cast in its weak form, then data is substituted directly into the equation, and parameters are estimated by solving a regression problem. A more detailed description of the method can be found in [26], and specifics about the implementation of WENDy in this study are given in Appendix A. Most importantly, this allows us to estimate parameters from systems in the form of (2) directly using WENDy.
@@ 3.2 Illustrating Examples
We are motivated by an example [1] where a weak-form input-output equation is used to estimate parameters in the blood-tissue diffusion model. Diffusion of drug between blood and tissue in the body is assumed to be proportional to the concentration gradient, and is represented by the following equation,
@@
where $x_{1}$ and $x_{2}$ are the concentrations of drug in the blood and tissue, respectively, $k_{12}$ is the diffusion rate from the blood into the tissue, $k_{21}$ is the diffusion rate from the tissue into the blood, and $V_{e}$ is the rate of drug decay in the blood. The input-output equation corresponding to (4) when only $x_{1}(t)$ is observed, can be written in the form of (1), and the weak-form input-output equation for the blood diffusion model is given below,
$\begin{split}-\int_{T_{1}}^{T_{2}}\ddot{\phi}x_{1}(t)\textsf{d}t=&w_{1}\int_{T_{1}}^{T_{2}}\phi\frac{x_{1}(t)}{x_{1}(t)+1}\textsf{d}t-w_{2}\int_{T_{1}}^{T_{2}}\dot{\phi}\frac{x_{1}(t)^{2}}{x_{1}(t)+1}\textsf{d}t+w_{3}\int_{T_{1}}^{T_{2}}\dot{\phi}\frac{1}{x_{1}(t)+1}\textsf{d}t,\end{split}$ (5)
where $w_{1}=k_{21}V_{e}$, $w_{2}=k_{12}+k_{21}$, and $w_{3}=k_{12}+k_{21}+V_{e}$.
@@
where $C$ is chosen so that $||\phi||_{2}=1$, $a$ is the radius of support, and $\eta$ is a shape parameter. For this example, $a=6$ and $\eta=9$.
We also consider Hartley modulating functions, which were applied to this weak-form system in [1]. Below we give the equation for a third order Hartley modulating function,
@@
where $C$ is chosen so that $||\phi||_{2}=1$, $a$ is the radius of support, and $\text{cas}(t)=\cos(t)+\sin(t)$. For this example, $a=8$.
@@
where $C$ is chosen so that $||\phi||_{2}=1$ and $a$ is the radius of support. For this example, $a=8$.
Figure 1 shows the drug diffusion system fit to only blood concentration using WENDy with a polynomial test function, and Figure 2 shows the average relative error using WENDy for various additive noise ratios for three different test functions. In comparison to [1], WENDy performs competitively, remaining under $50\%$ relative error with $80\%$ white noise in the observed variable when using a $12^{\text{th}}$ degree polynomial test function. This figure clearly demonstrates that the choice of test function greatly impacts the robustness of the parameter estimation method.
Figure 1: Using WENDy with equation (5) we recover parameters for the blood diffusion model (4) with only observations in the blood compartment ($x_{1}(t)$). The drug concentration dynamics in the blood compartment (red) and the tissue compartment (blue) are given for model (4) with parameters $k_{12}=5$, $k_{21}=1$, and $V_{e}=6$. We estimate the true parameters using WENDy from 400 observations of the blood compartment (black dots) with $e=5\%$ additive error ratio. Note that $e=5\%$ additive error ratio is equivalent to the white noise of $\sigma=25.06\%$ for this choice of model parameters.
Figure 2: Using WENDy with equation (5) is robust to noise. The average relative error for 1,000 simulated data sets with additive Gaussian error ratio $e\in[0\%,24.5\%]$ is given for choice of $C^{\infty}$ (see Eq. (6)), Hartley (see Eq. (7)), and Polynomial (see Eq. (8)) test functions. This is error ratio is equivalent to additive white noise of the level $\sigma\in[0,0.8].$
@@
The SIR model describes the spread of an infection in a population. Susceptible individuals ($S$) become infected ($I$) at a rate $\beta$ and recover ($R$) without chance of reinfection at a rate $\alpha$. Model equations are presented in the equation below,
@@
Note that, as expected, both parameters $\beta$ and $\alpha$ of model eq(9) are globally structurally identifiable. From equation (10) we can generate the weak form of model (9), SIR system for observed state $I$ is presented below,
@@
In this case we are unable to simplify the term $\int_{0}^{T}\phi\frac{\dot{I}^{2}}{I}\textsf{d}t$ using integration by parts. However, let the total population $N$ be constant, $N=S(t)+I(t)+R(t)$, and additionally, re-write $R$ in terms of only the observed variable $I$,
@@
Figure 3A shows the SIR model fit to only the infected population using WENDy.
Figure 3: Using WENDy with equations (13) we recover the transmission rate $\beta$ for the SIR model (9) with only observations in the infected compartment, even with large observational noise. The infected compartment dynamics are given for model (9) with N=10,000, $S_{0}=N-1$, $I_{0}=1$, $R_{0}=0$, $\beta=\frac{5.5}{N}$, $\gamma=5$. We estimate the true value of $\beta$ using WENDy from 31 observations of the infected compartment (black dots) and recover the original dynamics (red line) observations with $e=20\%$ additive error ratio (Right).
@@ 4 Examples of Practical Identifiability
In this section, we will describe the Practical identifiability of the weak form systems of the examples generated in Section 3.1. In Section 4.1, we describe the method of generating simulated datasets and parameter estimation. In Section 4.2, we interpret the practical identifiability of two biological examples.
@@ 4.1 Applying (e,q)-identifiability
To assess the practical identifiability of examples 3.1 and 3.2 we generated simulated data sets of size $N=1,000$ for choices of additive noise ratios between $e=[0\%,20\%]$ and $e=[0\%,80\%],$ respectively. We generated noisy observations using definition 3 and assuming a Gaussian additive error distribution. For each example, we assumed a limited number of data points; specifically, $40$ observations were considered for Example 3.1, and $31$ observations were considered for Example 3.2. Parameter estimates were generated using the WENDy method described in Section 3.1 and Appendix A.
Additionally, we compare the performance of WENDy to a traditional output error (OE) method. We implement the OE method in MATLAB using the function lsqnonlin with default settings. To perform the forward solves required by the OE method we use a variable-step, variable-order solver based on the numerical differentiation formulas (NDFs) of orders 1 to 5, as implemented in ode15s in Example 3.1 and a Dormand-Prince method as implemented in ode45 in Example 3.2.

### v3

@@ 3.1 Weak-form Input-output equations
In general, we convert from an equation in the form $\dot{U}=\Theta(U,p)$, with parameters $p$ and observed states, $y(t)=\Omega(U,p)$, to an input-output equation of the form $0=F(y,w)$, where the terms of $F(\cdot)$ are differential monomials and $w_{i}=f_{i}(p)$ are expressions of the parameters of the full ODE system. This process involves, first, ranking the model variables and their derivatives with the observed state variable at the lowest rank. Next, differential polynomials generated from the model equations are reduced to their characteristic set according to the new ranking using a differential elimination algorithm. After the characteristic set is obtained, the lowest ranking differential polynomial(s) will be the input-output equation(s).
@@
Let the right-hand side of the input-output equation can be written in the form
$F(y)=\sum_{i=0}^{n}\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}H_{i}(y){\color[rgb]{0,0,0}+\left(\sum_{i=0}^{n}\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}C_{i}(y)\right)w}.$ (1)
@@
Here, $H_{i}$ and $C_{i}$ are rational expressions of $y$. Then, it is clear that by convolving the input-output equation with a test function $\phi$ and applying integration by parts, we can generate a weak-form input-output equation of the form
$\color[rgb]{0,0,0}\sum_{i=0}^{n}\int_{0}^{T}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi(t))H_{i}(y)\textsf{d}t=\left(\sum_{i=0}^{n}\int_{0}^{T}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi(t))C_{i}(y)\textsf{d}t\right)w.$} (2)
However, not all nonlinear systems can be written in the form of equation (1). Instead, consider
$F(y)=\sum_{i=0}^{n}K_{i}(y)\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}H_{i}(y){\color[rgb]{0,0,0}+\left(\sum_{i=0}^{n}J_{i}(y)\frac{\textsf{d}^{i}}{\textsf{d}^{i}t}C_{i}(y)\right)w,}$ (3)
@@
where $K_{i}$ and $J_{i}$ are rational expression of $y$ and its derivatives $\{\dot{y}$, $\ddot{y},$...\}. In the case that the right-hand side of the input-output equation takes the form of equation (3), either substitutions of variables, estimation of higher order derivatives from data, or both will be required to compute terms that are not integrable by parts.
3.2 Weak form Estimation of Nonlinear Dynamics (WENDy)
@@
Recently, our group has developed the Weak form Estimation of Nonlinear Dynamics (WENDy) method for parameter estimation [28]. Briefly, in the WENDy method, a system is cast in its weak form, then data are substituted directly into the equation, and parameters are estimated by solving a regression problem. A more detailed description of the method can be found in [28]. Most importantly, this method allows us to estimate parameters from systems in the form of (2) directly using WENDy [^7] Note that in this formulation, we are limited to problems that are linear-in-parameters (not to be confused with linear problems). Recently, [49] has addressed problems that are nonlinear-in-parameters..
We begin by considering the ordinary differential equation (ODE) $\dot{U}=\Theta(U,p)$ with row vector of $d$ observed states
$y(t):=(y_{1}(t),...,y_{d}(t)).$
We assume the corresponding $d$-dimensional input-output equation $0=F(y,w)$ can be written in the form of (1). We let the superscript $\star$ notation denote quantities based on the true (noise-free and unknown) parameter values ($p^{\star}$). We assume that at each time point $t\in\mathbf{t}=(t_{0},\dots,t_{M})$ measurements of the system are observed with additive noise
$y(t)=y^{\star}(t)+\varepsilon(t)$ (4)
where each element of $\varepsilon(t)$ is i.i.d. $\mathcal{N}(0,\sigma^{2})$. Using bold-face variables to represent evaluation at the timegrid $\mathbf{t}$, our observations then consist of samples $\mathbf{y}:=\mathbf{y}^{\star}+\boldsymbol{\varepsilon}\in\mathbb{R}^{M+1\times d}$.
Our goal is to estimate the structurally identifiable parameter expressions $w^{\star}$ from the input-output equation representation of the ODE system. We start by considering the weak-form input-output equation
$\sum_{i=0}^{n}\int_{0}^{T}(-1)^{i+1}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi)H_{i}(y)\textsf{d}t=\left(\sum_{i=0}^{n}\int_{0}^{T}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi)C_{i}(y)\textsf{d}t\right)\textsf{mat}(w).$ (5)
where $\mathsf{mat}$” is the matricization operation.
Using a set of test functions $\{\phi_{k}\}_{i=1}^{K}$, (5) can be discretized at observed data $\mathbf{y}$, yielding approximated integrals
$\sum_{i=0}^{n}(-1)^{i+1}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi_{k})H_{i}(\mathbf{y})=\left(\sum_{i=0}^{n}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi_{k})C_{i}(\mathbf{y})\right)W.$ (6)
Here $\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi_{k}:=[\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{0})|...|\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{M})]\mathcal{Q}$ where $\mathcal{Q}$ is a diagonal matrix containing the Newton-Cotes quadrature weights. Let $\mathbf{C_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)C_{i}(\mathbf{y})$ and $\mathbf{H_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i+1}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)H_{i}(\mathbf{y})$. By minimizing the contribution of integration error to the residual, this discretization leads to the following corresponding linear system $\mathbf{G}\mathbf{w}^{\star}\approx\mathbf{b}$ [28], where
$\mathbf{G}$ $:=[\mathbb{I}_{d}\otimes\mathbf{C_{\phi,y}}],$
$\mathbf{b}$ $:=\mathsf{vec}(\mathbf{H_{\phi,y}}),$
from which we can obtain an estimate of $w^{\star}$.
As presented in [28], the WENDy algorithm addresses the Errors-In-Variable nature of the linearized problem using iterative reweighting to approximate the correct underlying covariance structure. This reweighting is dependent on the assumed noise structure of the observed data. For Gaussian additive noise, we apply the covariance correction using Algorithm 2 in [28]. Additionally, it has been shown in [49] that, through a strategic choice of transformations to the observations $\mathbf{y}$, it is possible to approximate the covariance structure for systems where observational noise is multiplicative and lognormal (i.e. $\mathbf{y}=\mathbf{y}^{\star}\varepsilon$ where $\log(\varepsilon(\mathbf{t}))\sim\mathcal{N}(0,\sigma^{2})$).
3.3 Structural Identifiablity of Weak-form Problem
Recall Definition 5 from Section 2.1, note that the terms of the input-output equation must be linearly independent on the problem’s domain in order to uniquely recover model parameters. It is also important to note that the conversion of an input-output equation in the form of equation (1) to a weak-form input-output equation in the form of equation (2) does not preserve the linear independence of terms for arbitrary choices of test function, $\phi(t)$. In particular, the integration operation does not preserve dependence relationships. Furthermore, the use of Definition 5 to check structural identifiability is equivalent to testing the membership in a differential field. An analogous problem in the weak-form would require the generation of an integro-differential field, and is an open field of research [50].
Alternatively, we propose verifying the recovery of the unique parameters, $w$, locally from the discretized problem given by equation (6). The WENDy regression problem has a unique solution when the matrix G is of full rank. In [28], this condition could be met for a choice of test functions such that $\text{rank}(\mathbf{\Phi})=K$ for a system without higher order derivatives of the observation states. We conjecture that for the case presented in Section 3.2, it is necessary to require $\text{rank}(\mathbf{\Phi})=\text{rank}(\dot{\mathbf{\Phi}})=\text{rank}(\mathbf{\frac{\textsf{{d}}^{i}}{\textsf{d}t^{i}}}\mathbf{\Phi})=K$ to obtain a unique solution to the WENDy regression problem. We compare the expected structural identifiability derived from the strong-form input-output equation to the local condition on the rank of matrix G for two examples in Appendix C.
Finally, it is essential to note that the conversion to a weak-form input-output equation entails the loss of information about any initial conditions. There are examples of biological systems where initial conditions are prerequisite to structural identifiability, for example, within-host models of virus dynamics [21, 37]. In this study, we consider examples of systems where structural identifiability is not dependent on initial conditions.
@@ 3.4 Illustrating Examples
We are motivated by an example [1] where a weak-form input-output equation is used to estimate parameters in the blood-tissue diffusion model. Diffusion of a drug between blood and tissue in the body is assumed to be proportional to the concentration gradient, and is represented by the following equation,
@@
where $x_{1}$ and $x_{2}$ are the concentrations of drug in the blood and tissue, respectively, $k_{12}$ is the diffusion rate from the blood into the tissue, $k_{21}$ is the diffusion rate from the tissue into the blood, and $V_{e}$ is the rate of drug decay in the blood. The input-output equation corresponding to (7) when only $x_{1}(t)$ is observed is given below,
$\begin{split}\ddot{x}_{1}(t)x_{1}(t)^{2}+2\ddot{x}_{1}(t)x_{1}(t)+\ddot{x}_{1}(t)+w_{1}\left(x_{1}(t)^{2}+x_{1}(t)\right)\\ +w_{2}\left(\dot{x}_{1}(t)x_{1}(t)^{2}+2\dot{x}_{1}(t)x_{1}(t)\right)+w_{3}\dot{x}_{1}(t)=0,\end{split}$ (8)
where $w_{1}=k_{21}V_{e}$, $w_{2}=k_{12}+k_{21}$, and $w_{3}=k_{12}+k_{21}+V_{e}$. We confirm that equation (8) is globally structurally identifiable using available software in Appendix C.
Equation (8) can be written in the form of (1) and the weak-form input-output equation for the blood diffusion model is given below,
$\begin{split}-\int_{T_{1}}^{T_{2}}\ddot{\phi}x_{1}(t)\textsf{d}t=&w_{1}\int_{T_{1}}^{T_{2}}\phi\frac{x_{1}(t)}{x_{1}(t)+1}\textsf{d}t-w_{2}\int_{T_{1}}^{T_{2}}\dot{\phi}\frac{x_{1}(t)^{2}}{x_{1}(t)+1}\textsf{d}t+w_{3}\int_{T_{1}}^{T_{2}}\dot{\phi}\frac{1}{x_{1}(t)+1}\textsf{d}t.\end{split}$ (9)
@@
where $C$ is chosen so that $\phi||_{2}=1$, $a$ is the radius of support, and $\eta$ is a shape parameter. For this example, $a=0.6$ and $\eta=9$.
We also consider Hartley modulating functions, which were applied to this weak-form system in [1]. Below, we give the equation for a third-order Hartley modulating function,
@@
where $C$ is chosen so that $\phi||_{2}=1$, $a$ is the radius of support, and $\text{cas}(t)=\cos(t)+\sin(t)$. For this example, $a=0.8$.
@@
where $C$ is chosen so that $\phi||_{2}=1$ and $a$ is the radius of support. For this example, $a=0.52$.
Figure 1 shows the drug diffusion system fit to only blood concentrations using WENDy with a polynomial test function, and Figure 2 shows the average relative error using WENDy for various additive noise ratios for three different test functions. In comparison to [1], WENDy performs competitively, remaining under $50\%$ for the relative error with $80\%$ white noise in the observed variable when using a $12^{\text{th}}$ degree polynomial test function. This figure clearly demonstrates that the choice of test function greatly impacts the robustness of the parameter estimation method. For further details on the choice of test functions and their respective radii of support, see Appendix A.
Figure 1: Using WENDy with equation (9) we recover parameters for the blood diffusion model (7) with only observations in the blood compartment ($x_{1}(t)$). The drug concentration dynamics in the blood compartment (red) and the tissue compartment (blue) are given for model (7) with parameters $k_{12}=5$, $k_{21}=1$, and $V_{e}=6$. These parameters are chosen to allow comparison to the example in [1]. The plot depicts the relative error in estimating the true parameters using WENDy from 400 observations of the blood compartment (black dots) with $e=5\%$ additive observation error ratio. Note that $e=5\%$ additive observation error ratio is equivalent to the white noise of $\sigma=25.06\%$ for this choice of model parameters.
Figure 2: For using WENDy to estimate the parameters in (9), this figure depicts the relative parameter error vs. observation error ratio $e$. The different curves represent using WENDy with different test functions ($C^{\infty}$ from (10), Hartley from (11), and Polynomial from (12)). When using either $C^{\infty}$ or polynomial test functions (as is standard in WENDy), the relative error remains below $25\%$ for observations with up to a $10\%$ additive error ratio, and for polynomial test functions, the relative error remains below $50\%$ for observations with up to a $24.5\%$ additive error ratio. Note that each value on these curves represents the average relative error for 1,000 simulated datasets with additive Gaussian error ratio $e\in[0\%,24.5\%]$. Also note that the error ratio is equivalent to additive white noise of the level $\sigma\in[0,0.8].$
@@
The SIR model describes the spread of an infection in a population with no birth, no death, nor immigration. Susceptible individuals ($S$) become infected ($I$) at a rate $\beta$ and recover ($R$) without chance of reinfection at a rate $\alpha$. Model equations are presented in the equation below,
@@
Note that, as expected, both parameters $\beta$ and $\alpha$ of model (13) are globally structurally identifiable. We confirm that equation (14) is globally structurally identifiable using available software in Appendix C.
From equation (14) we can generate the weak form of model (13), SIR system for observed state $I$ is presented below,
@@
In this case, we are unable to simplify the term $\int_{0}^{T}\phi\frac{\dot{I}^{2}}{I}\textsf{d}t$ using integration by parts. However, let the total population $N$ be a known constant, $N=S(t)+I(t)+R(t)$, and additionally, re-write $R$ in terms of only the observed variable $I$,
@@
We verify the structural identifiability of equation (14) and the local recovery of parameters in the weak form Equation (17) in Appendix C. Figure 3A shows the SIR model fit to only the infected population using WENDy.
Figure 3: Using WENDy with equations (17) we recover the transmission rate $\beta$ for the SIR model (13) with only observations in the infected compartment, even with large observational noise. The infected compartment dynamics are given for model (13) with N=10,000, $S_{0}=N-1$, $I_{0}=1$, $R_{0}=0$, $\beta=\frac{5.5}{N}$, $\gamma=5$. These parameter values correspond to an infection with a 5-day infectious period and a basic reproduction number $\mathcal{R}_{0}=1.1$, which could reasonably represent a seasonal influenza outbreak [51]. We estimate the true value of $\beta$ using WENDy from 31 observations of the infected compartment (black dots) and recover the original dynamics (red line) observations with $e=20\%$ additive observation error ratio (Right).
@@ 4 Examples of Practical Identifiability
In this section, we will describe the Practical identifiability of the weak form systems of the examples generated in Section 3.1. In Section 4.1, we describe the method of generating simulated datasets. In Section 4.2, we interpret the practical identifiability of two biological examples.
@@ 4.1 Applying (e,q)-identifiability
To assess the practical identifiability of examples 3.1 and 3.2 we generated $D=1,000$ simulated datasets for choices of observational error ratios between $e=[0\%,20\%]$. We generated noisy observations using Definition 3 and assuming either a Gaussian additive error distribution or a multiplicative log-normal distribution. For each example, we assumed a limited number of data points; specifically, $40$ observations were considered for Example 3.1, and $31$ observations were considered for Example 3.2. Parameter estimates were generated using the WENDy method described in Section 3.2 with test functions given in Appendix A.
Additionally, we compare the performance of WENDy to a traditional output error (OE) method. We implement the OE method in MATLAB using the function lsqnonlin with default settings. To perform the forward solves required by the OE method, we use a variable-step, variable-order solver based on the numerical differentiation formulas (NDFs) of orders 1 to 5, as implemented in ode15s in Example 3.1 and a Dormand-Prince method as implemented in ode45 in Example 3.2.